OpenStudy (just_one_last_goodbye):

If cos(53) = 0.6 what is sine of it's complement?

OpenStudy (bookworm14):

@mathstudent55 @ikram002p @ParthKohli

OpenStudy (wendy.ivette11714):

Do you have any idea?

OpenStudy (just_one_last_goodbye):

well not really ._.

OpenStudy (wendy.ivette11714):

Eh same I suck at math

OpenStudy (just_one_last_goodbye):

should I post another answer?

OpenStudy (wendy.ivette11714):

Answer or question?

OpenStudy (just_one_last_goodbye):

question http://i.imgur.com/HyFguVj.png

OpenStudy (wendy.ivette11714):

Have any guess?

OpenStudy (just_one_last_goodbye):

Jiggly :D help me pwease ;-;

jigglypuff314 (jigglypuff314):

*looks at the question* *looks at the screenshot* do you mean that x = 53 and adjacent over hypotenuse = 3 over 5 ?

OpenStudy (just_one_last_goodbye):

this is what it says http://i.imgur.com/pkOFODK.png

OpenStudy (just_one_last_goodbye):

thats for the first one

OpenStudy (just_one_last_goodbye):

second is just to find x.... but idk what all those other 3 are for

jigglypuff314 (jigglypuff314):

"The cosine of an acute angle is equal to the sine of its complement." http://www.regentsprep.org/regents/math/algtrig/att6/cofunctions.htm

OpenStudy (just_one_last_goodbye):

so... it's 0.6?

jigglypuff314 (jigglypuff314):

I think so ^_^"

OpenStudy (just_one_last_goodbye):

thanks :D what about the second?

jigglypuff314 (jigglypuff314):

I guess you just use variables... csc is 1/sin = 1/(opp/hyp) = hyp/opp sec is 1/cos cot is 1/tan

rvc (rvc):

did u mean sin(90-x)=cosx jiggly?

jigglypuff314 (jigglypuff314):

I have no clue :P @rvc you would know better than me, what do you think?

OpenStudy (just_one_last_goodbye):

thank so much Jiggly <3

OpenStudy (just_one_last_goodbye):

meaning?

OpenStudy (anonymous):

\(\alpha\) is alpha \(\beta\) is beta \(\gamma\) is gamma These are what we use for angles in trignometry.

OpenStudy (just_one_last_goodbye):

so whats csc = ?

OpenStudy (anonymous):

csc is 1/sin

OpenStudy (anonymous):